OpenStudy (zmudz):

Find the maximum of \(2x + 2\sqrt{x(1-x)}\) over \(0 \leq x \leq 1.\)

OpenStudy (thomas5267):

Olympiad or calculus question?

OpenStudy (thomas5267):

\[ f(x)=2x+2\sqrt{x(1-x)}\\ f(x)=2x+2\sqrt{\left(-x^2+x\right)}\\ f'(x)=2+\left(-x^2+x\right)^{-\frac{1}{2}}(-2x+1) \] \[ \begin{align*} f'(x)&=0\\ 2+\left(-x^2+x\right)^{-\frac{1}{2}}(-2x+1)&=0\\ \left(-x^2+x\right)^{-\frac{1}{2}}(-2x+1)&=-2\\ -2x+1&=-2\left(-x^2+x\right)^{\frac{1}{2}}\\ 2x-1&=2\left(-x^2+x\right)^{\frac{1}{2}}\\ 4x^2-4x+1&=4\left(-x^2+x\right)\\ 8x^2-8x+1&=0\\ x=\frac{8+\sqrt{32}}{16}&\text{ or }x=\frac{8-\sqrt{32}}{16}\\ x=\frac{2+\sqrt{2}}{4}&\text{ or }x=\frac{2-\sqrt{2}}{4}\\ \end{align*} \] \[ \begin{align*} f\left(\frac{2+\sqrt{2}}{4}\right)&=\frac{2+\sqrt{2}}{2}+2\sqrt{-\left(\frac{1}{4}+\frac{\sqrt{2}}{4}+\frac{1}{8}\right)+\frac{2+\sqrt{2}}{4}}\\ &=\frac{2+\sqrt{2}}{2}+2\sqrt{\frac{1}{8}}\\ &=\frac{2+\sqrt{2}}{2}+\frac{\sqrt{2}}{2}\\ &=1+\sqrt{2}\\\\ f\left(\frac{2-\sqrt{2}}{4}\right)&=\frac{2-\sqrt{2}}{2}+2\sqrt{-\left(\frac{1}{4}-\frac{\sqrt{2}}{4}+\frac{1}{8}\right)+\frac{2-\sqrt{2}}{4}}\\ &=\frac{2-\sqrt{2}}{2}+2\sqrt{\frac{1}{8}}\\ &=\frac{2-\sqrt{2}}{2}+\frac{\sqrt{2}}{2}\\ &=1-\sqrt{2} \end{align*}\\ f\left(\frac{2+\sqrt{2}}{4}\right)=1+\sqrt{2}\text{ is the maximum.} \]

OpenStudy (thomas5267):

Took me 40+ minutes since I am really tired.

OpenStudy (zmudz):

it was a precalc question